y=(x^3 -3x^2)/(2x-1)
Answer by ewatrrr(24785)

 
Hi,
y=(x^3 -3x^2)/(2x-1) = .5x^2 - 1.25x - .625 R = -.625
Found by doing routine long division
2x -1|x^3 - 3x^2